This problem comes from the Accuplacer practice test.
Sarah needs to make a cake and some cookies. The cake requires 3/8 cup of sugar and the cookies require 3/5 cup of sugar. Sarah has 15/16 cups of sugar. Does she have enough sugar, or how much more does she need?

She has enough sugar.
She needs 1/8 of a cup of sugar.
She needs 3/80 of a cup of sugar.
She needs 4/19 of a cup of sugar.
She needs 1/9 of a cup of sugar.

I can get as far as: 3/8 + 3/5 -> 15/40 + 24/40 = 39/40. I know the answer is 3/80C. of Sugar, but I can't figure out how they worked it out. I've tried several variations, but cannot = 3/80.

Answer by stanbon(75887)

3/8 cup of sugar and the cookies require 3/5 cup of sugar. Sarah has 15/16 cups of sugar.
3/8 +3/5 = (15+24)/40= 39/40
15/16R39/40
Which is greater? 15/16 or 39/40
Crossmultiply to get:
15*39R16*40
585R640
As you see 585<640
Therefore 15/16 is less that 39/40
Sarah does not have enough sugar.
How much more does she need?
39/40-15/16
=78/80- 75/80
= 3/80 of a cup of sugar
Cheers,
Stan H.